04 March 2014

More clouds

Today continues the frustratingly cloudy weather we've been having for a couple of days. Temperatures remain relatively warm. Eureka was -29°C today, with occasional light snow. Outside PEARL it was -26°C.

Before the Ridge Lab's daily commute after breakfast, the Eureka Weather Station manager, John, let Emily, Dan, and Zahra ride along with him in the (very cool) Environment Canada trax truck to Upper Paradise. John was asked to fix something for the Canadian Forces. Zahra and Dan planned to then catch a ride the rest of the way in the CANDAC truck, but the road wasn't in good shape. Winds and snow had created snow drifts overnight. John offered to continue on to the Ridge Lab once he was done to help the CANDAC truck pass through the drifts easier. The team was thankful for the offer. It made the ride much easier, and it gave Emily a chance to briefly visit the Ridge Lab before she heads home in a couple of days.

The Ridge Lab's IR lab was quiet again today. Dan and Joseph couldn't take measurements with the Bruker and PARIS due to clouds.

Over in the UV-Vis Lab, Sophie noticed the UT-GBS's spectra were saturated this morning. She put a diffuser inside the telescope on the optical fiber spot end. The spectra look good now. Sophie needs to record a full day of measurements to do a quick QDOAS analysis and check what impact the new diffuser is having on the spectra. She has also installed temperature sensors on UT-GBS Triax spectrometer, its CCD, and telescope.

Today, Zahra and Tom tried to align the SPS and MAESTRO suntracker with the Sun. This was done manually, but the tracker still could not find its zero. After trying many different approaches, they got the tracker working well. Angles between where the tracker was pointing and the Sun were calculated and converted into the number of steps the tracker needed to rotate to find the Sun. These corrections were added to the file the tracker reads its position from, and the tracker azimuth was aligned. The elevation was aligned manually. Tom and Zahra will run another test tomorrow to ensure the alignment is robust.

SPS and MAESTRO did not perform measurements during the alignment procedures. Before the team left the Ridge Lab for the day, Tom and Zahra started the instruments again. SPS and MAESTRO are again making nominal measurements.

The Brewer had caused Tom and Zahra concern yesterday. It had not been working since installed on the Ridge Lab roof. They had been concerned there might be a problem with its firmware. Before diving into firmware changes, Tom decided to use the cable he used to perform tests with inside the UV-Vis Lab. Tom and Zahra ran the cable down from the roof to the lab and connected the Brewer. It worked! The problem had been two broken cables. That was a relief. Tom tried to sight the Brewer, but there was too little sunlight. He will have to sight it tomorrow.

Down at 0PAL, the CRL ran for 12 hours in all UV and Vis channels overnight until this morning. Emily and Sham performed depolarized calibrations for 2 hours this afternoon. They resumed regular measurements just before suppertime, and are set to continue all night. Time was also spent installing software on Sham's computer for data acquisition and analysis, as well as strategizing for water vapour channel tests which Sham plans to run soon.

Today's ACE ozonesonde flight occurred at 6:15 PM (local time) and reached an altitude of 32 253 meters (7.6 hPa).

In the evening, Tom gave the Eureka Weather Station and the PEARL team a talk on his trip to the South Pole and the history of the Eureka Weather Station. After that, the team watched a documentary on Nunavut, Arctic Defenders .

- Dan
Back to Daily Reports
Back to Home